# Using Symbolic Regression to Emulate the Radial Fourier Transform of the Sérsic profile for Fast, Accurate and Differentiable Galaxy Profile Fitting

### v1

@@ 1 Introduction
Galaxy profile fitting describes a forward modeling approach to measuring the structural parameters and photometry of galaxies from their images. In many cases the galaxy is assumed to be axisymmetric, with a parametric form of the radial surface brightness profile. The best fit parameters are used to investigate the shape of the surface brightness profile, size, and total flux of the galaxy. By directly modeling the image, one is able to account for the effect of the point spread function, or PSF. Profile modeling has thus become a ubiquitous tool for modern extragalactic astronomy for investigating the morphological properties of galaxies (e.g. Lange et al., 2015; Mowla et al., 2019; Kawinwanichakij et al., 2021; Ward et al., 2024; Allen et al., 2024). For large surveys it is used to obtain accurate photometry of extended sources. (Bernardi et al., 2013) Such modeling is also necessary to accurately model sources in crowded fields and across instruments with vastly different depths and spatial resolutions. (Dey et al., 2019; Weaver et al., 2022)
@@
$I(R)\propto F_{\rm total}\exp\left[\left(\frac{R}{R_{\rm eff}}\right)^{1/n}\right]$ (1)
Here, $F_{\rm total}$ is the total flux of the galaxy, $n$ is the Sérsic index and describes the shape of the profile, and $R_{\rm eff}$ is the effective — or half-light — radius, which describes the spatial scale of the galaxy. These three parameters, along with the axis ratio, position angle and central position, are used to parameterize the light profile of galaxies. The introduction of the index, $n$, consolidated the fixed index profiles such as the de Vaucouleurs (de Vaucouleurs, 1948) and exponential into a single generalized parametrization. Given this additional freedom the Sérsic profile is able to provide a reasonable approximation to the surface brightness profiles of almost all galaxies. There are numerous software packages and methods that perform Sérsic profile fitting, each with their own aims and strengths. Some of the most popular examples are: galfit (Peng et al., 2002), imfit (Erwin, 2015), profit (Robotham et al., 2017), astrophot (Stone et al., 2023), pysersic (Pasha & Miller, 2023) and forcepho (Baldwin et al., 2024).
@@
A promising alternative method is to render the galaxy in Fourier space; with this method, similar accuracy in the central region can be achieved without the need for oversampling (Rowe et al., 2015; Lang, 2020). This provides a more natural way to render the galaxy “above the atmosphere,” i.e., the intrinsic galaxy image before including the effects of the PSF and pixelization. Furthermore, a Fourier transform is required to convolve a pixel-rendered image with a PSF regardless; rendering directly in Fourier space negates the need for this initial computation, improving performance. Unfortunately, despite the theoretical benefits of a Fourier representation, the Sérsic function has no closed-form Fourier transform.
A popular workaround, originally suggested in Hogg & Lang (2013), is to approximate the Sérsic profile using a mixture of Gaussians model. Gaussians are particularly nice functions to work with, especially when using Fourier transforms, as the transform of a Gaussian is also Gaussian. Gaussian decompositions of Sérsic profiles can either be pre-computed and stored in a lookup table (Rowe et al., 2015; Lang et al., 2016) or calculated on the fly, such as using the algorithm presented in  Shajib (2019). Additionally, if the PSF is also modeled using Gaussians, then the convolution can be performed analytically. The Mixture of Gaussians representation has been implemented in packages for Sérsic profile fitting such as tractor (Lang et al., 2016), pysersic (Pasha & Miller, 2023) and forcepho (Baldwin et al., 2024). Unfortunately this solution is not a panacea — in order to accurately reproduce the Sérsic profiles, a large number of Gaussians ($\gtrsim 10$) must be employed. Each Gaussian is rendered independently, thus (again) increasing the computational cost.
@@
In this paper, we tackle this longstanding problem using a novel approach. Our first insight is recognizing that although the Fourier transform of the Sérsic profile is intractable to calculate analytically, it is possible to calculate numerically. Thus one can easily build up a library of curves that could be used to train an emulator to mimic its behavior. Our second insight is to employ symbolic regression to emulate the Fourier transforms of the Sérsic profile. Symbolic regression is a supervised learning task which aims to search a wide space of functional forms to find the best fit to a given set of data. (Angelis et al., 2023) This is usually achieved using a genetic programming approach which begins with a set of simple expressions and combines them to generate more complex equations. For this study we will use symbolic regression to mimic the Fourier transform of the Sérsic profile to build an emulator based on simple, computationally efficient and differentiable building blocks. This last point is particularly salient given the rise of gradient based optimization and sampling techniques being implemented for profile fitting (Stone et al., 2023; Pasha & Miller, 2023; Chen et al., 2024).
The paper is organized as follows: First we calculate the radially symmetric Fourier transform numerically in Section 2. We then emulate these curves utilizing the pysr package (Cranmer, 2023) to perform symbolic regression in Section 3. In Section 4 we implement a rendering method based on this emulator in the pysersic package to ensure its accuracy for profile fitting We summarize our approach and discuss caveats and future directions in Section 5.
@@ 2 The Radially Symmetric Fourier Transform of the Sérsic Profile
We use the python package hankel (Murray & Poulin, 2019), to numerically calculate the radial Fourier transform. By using the numerical integration method for Bessel functions derived in Ogata (2005), this package is able to quickly and accurately calculate the Hankel transform for any given function. For all of the calculations performed in this paper, we use a Sérsic profile that is normalized with a total luminosity and effective radius equal to unity. For all calculations performed in this study using hankel, the quadrature bin width, $h$, which controls the accuracy of the integrations is set to $10^{-4}$. Correspondingly, the number of quadrature bins, $N$, we set to $\pi/h$ as suggested in the documentation. This level of accuracy is likely unnecessary, as the reported uncertainties are often lower than machine precision, $10^{-16}$, however all of the transforms needed for this study are able to be calculated within a few minutes.
@@ 3.1 Fitting Procedure
Figure 2: Showcasing the results of the symbolic regression fitting process comparing execution time and two measures of accuracy. The execution time is measured for rendering a model image, calculating a $\chi^{2}$ log-likelihood and calculating the gradients of the input parameters. We compare this against the final loss from the symbolic regression. Each point is colored by its complexity. We find a negative correlation between execution time and accuracy; more complex equations tend to be more accurate but take longer to compute. However above a complexity of 30 there is a plateau in accuracy. In order to balance computational efficiency and accuracy we select an equation near the bottom of this plateau, denoted by $\tilde{\mathcal{F}}_{r}$ and shown in Eqn. 6
@@
where $G(k,n)$ is an arbitrary function that will be searched for using symbolic regression. This constraint is imposed using the template functionality in pysr [^1] Introduced in version 1.0. This form, essentially a Sigmoid function, enforces the boundedness observed in Fig. 1. Initially in our testing we initially did not enforce any structure but found many of the resulting equations eventually converged to a similar form as presented in Eqn. 3. Enforcing this structure a-priori allows for more efficient convergence that results in more accurate final equations.
@@
We construct our training set by numerically randomly sampling 10,000 values of $k,n$ and using the hankel python package to calculate $\mathcal{F}_{r}$ $(k,n)$. In this calculation, the flux and $R_{\rm eff}$ of the profile are set to unity. The resulting equation can be trivially rescaled to any value and this normalization allows us to focus the variation in $\tilde{\mathcal{F}_{r}}$ $(k,n)$ with $k$ and $n$. For $n$ we randomly sample values between 0.4 and 6.5, using a uniform distribution. Our target range for $n$ is 0.560.5-6 and by extending the window slightly for the training set we aim to limit the decrease in accuracy near the edges target range. We sample the set of $k$ values from an exponential distribution with a rate of $1/15$. We find that this roughly mimics the distribution of $k$ values when a 200 pixel by 200 pixel image is Fourier transformed.
@@
In general, we use most of the default options provided by pysr. We allow the following basic binary operators: +,,,+,-,*,, and the following unary operators: $\exp,\log,{\rm square},{\rm sqrt}$, inverse and cube. Setting the inverse as a unary operator, instead of the binary division operator, allows us to better control the constraints to avoid numerical issues. We set up a series of constraints on these operators that attempt to avoid circularly nested functions, i.e. a sqrt directly inside a square, or a log directly inside an exponential. Additionally we do not allow any ++ or - operators inside the inverse to avoid division by 0. We use a L2, or mean square error, loss function and a maximum equation complexity of 50. We perform the symbolic search for $2.5\times 10^{6}$ iterations, with the results summarized in Figure 2.
@@ 3.2 Selecting an Equation
For many applications, the goal of symbolic regression is to find a simple or interpretable equation. However, our aim is to find an accurate and computationally efficient equation, with little emphasis on readability or simplicity. To understand the properties of the final population of equations in this context, Figure  2 compares the computational efficiency against the accuracy of the final pool of equations. For computational efficiency we measure the execution time to render an 150 x 150 pixel image for a given set of parameters and calculate the $\chi^{2}$ log likelihood with respect to a reference image. This is implemented in jax, same as pysersic, which allows for auto-differentiation capabilities. In this timing we also include the calculation of gradients of the log likelihood. This setup attempts to mimic the calculations used for sampling and optimization when these emulators are implemented into pysersic. This timing is repeated 500 times for each equation and we report the median.
@@
Figure 3: Injection recovery tests to ensure our chosen Sérsic Fourier Emulator, Eqn. 6 accurately for profile fitting. We render a series of profiles using a heavily over-sampled pixel space renderer as the ‘truth‘ and fit each using a renderer based on the Fourier Emulator. The input values are compared to those recovered for $r_{\rm eff}$ and the index $n$. We find great agreement, with the average residual $0.5\%$ and the scatter $2\%$, calculated using the biweight location and scale respectively (Beers et al., 1990), for both parameters. These results indicate that the rendering method utilizing the Sérsic Fourier emulator can be used for profile fitting with minimal loss in accuracy.
@@
In general we find there is an inverse correlation between accuracy and computational time. The more complex equations are generally more accurate but require more calculations and therefore have longer execution times. We find there is a plateau in this relationship. Above a complexity of about 30 the execution time continues to increase without any significant gains in the accuracy. In order to maximize both computational efficiency and accuracy we aim to select a point near the beginning of this plateau. Specifically we select the equation with the lowest execution time that has L2 loss $2\times 10^{-6}$. This yields the equation with a complexity of 39, which is displayed below, in Equation. 6.
@@
This equation is an outlier in the plane of execution time and loss that makes it particularly suited for our use case. Compared to equations of similar complexity in the final pool there are very few instances of the $k$ evaluation within exponential or square root evaluations. When rendering a profile there number of values of $k$ that need to be calculated scales with the number of pixels and these are computationally expensive operations. Given there is only one value of $n$ to evaluate for each profile the prevalence of this variable has less of an effect on the overall execution time. For the remainder of this paper we will use this equation used to emulate the radial Fourier transform of the Sérsic profile.

### v2

@@ 1 Introduction
Galaxy profile fitting describes a forward modeling approach to measuring the structural parameters and photometry of galaxies from their images. In many cases the galaxy is assumed to be axisymmetric, with a parametric form of the radial surface brightness profile. The best fit parameters are used to investigate the shape of the surface brightness profile, size, and total flux of the galaxy. By directly modeling the image, one is able to account for the effect of the point spread function, or PSF. Profile modeling has thus become a ubiquitous tool for modern extragalactic astronomy for investigating the morphological properties of galaxies (e.g. Lange et al., 2015; Mowla et al., 2019; Kawinwanichakij et al., 2021; Ward et al., 2024; Allen et al., 2024). For large surveys it is used to obtain accurate photometry of extended sources (Bernardi et al., 2013). Such modeling is also necessary to accurately model sources in crowded fields and across instruments with vastly different depths and spatial resolutions (Dey et al., 2019; Weaver et al., 2022).
@@
$I(R)\propto F_{\rm total}\exp\left[-\left(\frac{R}{R_{\rm eff}}\right)^{1/n}\right]$ (1)
Here, $F_{\rm total}$ is the total flux of the galaxy, $n$ is the Sérsic index and describes the shape of the profile, and $R_{\rm eff}$ is the effective — or half-light — radius, which describes the spatial scale of the galaxy. These three parameters, along with the axis ratio, position angle and central position, are used to parameterize the light profile of galaxies. The introduction of the index, $n$, consolidated the fixed index profiles such as the de Vaucouleurs (de Vaucouleurs, 1948) and exponential into a single generalized parametrization. Given this additional freedom the Sérsic profile is able to provide a reasonable approximation to the surface brightness profiles of almost all galaxies. It is also possible to create non-axisymmetric models from this base profile by adding Fourier modes or warps (Peng et al., 2002; Stone et al., 2023). There are numerous software packages and methods that perform Sérsic profile fitting, each with their own aims and strengths. Some of the most popular examples are: galfit (Peng et al., 2002), imfit (Erwin, 2015), profit (Robotham et al., 2017), astrophot (Stone et al., 2023), pysersic (Pasha and Miller, 2023) and forcepho (Baldwin et al., 2024).
@@
A promising alternative method is to render the galaxy in Fourier space; with this method, similar accuracy in the central region can be achieved without the need for oversampling (Rowe et al., 2015; Lang, 2020). This provides a more natural way to render the galaxy “above the atmosphere,” i.e., the intrinsic galaxy image before including the effects of the PSF and pixelization. Furthermore, a Fourier transform is required to convolve a pixel-rendered image with a PSF regardless; rendering directly in Fourier space negates the need for this initial computation, improving performance. Unfortunately, despite the theoretical benefits of a Fourier representation, the Sérsic function has no general closed-form Fourier transform [^1] At fixed indices such as $n=0.5$, representing a Gaussian, and $n=1$ representing an exponential the Fourier transforms are known analytically.
A popular workaround, originally suggested in Hogg and Lang (2013), is to approximate the Sérsic profile using a mixture of Gaussians model. Gaussians are particularly nice functions to work with, especially when using Fourier transforms, as the transform of a Gaussian is also Gaussian. Gaussian decompositions of Sérsic profiles can either be pre-computed and stored in a lookup table (Rowe et al., 2015; Lang et al., 2016) or calculated on the fly, such as using the algorithm presented in Shajib (2019). Additionally, if the PSF is also modeled using Gaussians, then the convolution can be performed analytically. The Mixture of Gaussians representation has been implemented in packages for Sérsic profile fitting such as tractor (Lang et al., 2016), pysersic (Pasha and Miller, 2023) and forcepho (Baldwin et al., 2024). Unfortunately this solution is not a panacea — in order to accurately reproduce the Sérsic profiles, a large number of Gaussians ($\gtrsim 10$) must be employed. Each Gaussian is rendered independently, thus (again) increasing the computational cost.
@@
In this paper, we tackle this longstanding problem using a novel approach. Our first insight is recognizing that although the Fourier transform of the Sérsic profile is intractable to calculate analytically, it is possible to calculate numerically. Thus one can easily build up a library of curves that could be used to train an emulator to mimic its behavior. Our second insight is to employ symbolic regression to emulate the Fourier transforms of the Sérsic profile. Symbolic regression is a supervised learning task which aims to search a wide space of functional forms to find the best fit to a given set of data (Angelis et al., 2023). This is usually achieved using a genetic programming approach which begins with a set of simple expressions and combines them to generate more complex equations. For this study we will use symbolic regression to mimic the Fourier transform of the Sérsic profile to build an emulator based on simple, computationally efficient and differentiable building blocks. This last point is particularly salient given the rise of gradient based optimization and sampling techniques being implemented for profile fitting (Stone et al., 2023; Pasha and Miller, 2023; Chen et al., 2024).
The paper is organized as follows: First we calculate the radially symmetric Fourier transform numerically in Section 2. We then emulate these curves utilizing the pysr package (Cranmer, 2023) to perform symbolic regression in Section 3. In Section 4 we implement a rendering method based on this emulator in the pysersic package to ensure its accuracy for profile fitting. We summarize our approach and discuss caveats and future directions in Section 5.
@@ 2 The Radially Symmetric Fourier Transform of the Sérsic Profile
We use the python package hankel (Murray and Poulin, 2019), to numerically calculate the radial Fourier transform. By using the numerical integration method for Bessel functions derived in Ogata (2005), this package is able to quickly and accurately calculate the Hankel transform for any given function. For all of the calculations performed in this paper, we use a Sérsic profile that is normalized with a total luminosity and effective radius equal to unity. For all calculations performed in this study using hankel, the quadrature bin width, $h$, which controls the accuracy of the integrations is set to $10^{-4}$. Correspondingly, the number of quadrature bins, $N$, we set to $\pi/h$ as suggested in the documentation. This level of accuracy is likely unnecessary, as the reported uncertainties are often lower than machine precision, $10^{-16}$, however all of the transforms needed for this study are able to be calculated within a few minutes.
@@ 3.1 Fitting Procedure
Figure 2: Showcasing the results of the symbolic regression fitting process comparing execution time and accuracy. The execution time is measured for rendering a model image, calculating a $\chi^{2}$ log-likelihood and calculating the gradients of the input parameters. The accuracy is calculated using the L2 loss, or mean-squared error, of a validation set of $k,n$-$\mathcal{F}_{r}$ pairs. Each point is colored by its complexity. We find a negative correlation between execution time and accuracy; more complex equations tend to be more accurate but take longer to compute. However above a complexity of 30 there is a plateau in accuracy. In order to balance computational efficiency and accuracy we select an equation near the bottom of this plateau, denoted by $\tilde{\mathcal{F}}_{r}$ and shown in Eqn. 6. For comparison we note that the default hybrid renderer in pysersic has an execution time of 1.66 ms, more than twice as slow as the most complex equations presented here and three times slower than our selected equation.
@@
where $G(k,n)$ is an arbitrary function that will be searched for using symbolic regression. This constraint is imposed using the template functionality in pysr [^2] Introduced in version 1.0. This form is a Sigmoid function with the opposite sign in the exponential to match the limits of the limits of $\mathcal{F}_{r}$and enforces the boundedness observed in Fig. 1. Initially in our testing, we did not enforce any structure but found many of the resulting equations eventually converged to a similar form as presented in Eqn. 3. Enforcing this structure a-priori allows for more efficient convergence that results in more accurate final equations. Another option to enforce this boundedness would instead forgo using the template and perform the inverse sigmoid transform before fitting, i.e. performing symbolic regression on $\log\left(\frac{1}{1-\mathcal{F}_{r}}\right)$. However this, in essence, optimizes for fractional uncertainty on $\mathcal{F}_{r}$, and places too much emphasis on the high and low values of $k$. Once the sigmoid function has been applied, minor differences in this regime do not alter the behavior of the resulting equation much. Using the template approach allows us to optimize for absolute uncertainty on $\mathcal{F}_{r}$ rather than fractional uncertainty.
@@
We construct our training set by numerically randomly sampling 10,000 values of $k,n$ and using the hankel python package to calculate $\mathcal{F}_{r}$ $(k,n)$. In this calculation, the flux and $R_{\rm eff}$ of the profile are set to unity. The resulting equation can be trivially rescaled to any value and this normalization allows us to focus the variation in $\tilde{\mathcal{F}_{r}}$ $(k,n)$ with $k$ and $n$. For $n$ we randomly sample values between 0.4 and 6.5, using a uniform distribution. Our target range for $n$ is 0.560.5-6, encompassing the majority of real galaxies (Lange et al., 2015; Mowla et al., 2019) and extend the window slightly for the training set to limit the decrease in accuracy near the edges target range. We sample the set of $k$ values from an exponential distribution with a rate of $1/15$. We find that this roughly mimics the distribution of $k$ values when a 200 pixel by 200 pixel image is Fourier transformed. This is around the typical size of cutout used for galaxy image modeling, which is often between 100-200 pixels. (van der Wel et al., 2012; Mowla et al., 2019).
@@
We allow the following basic binary operators: +,,,+,-,*,, and the following unary operators: $\exp,\log,{\rm square},{\rm sqrt}$, inverse and cube. Setting the inverse as a unary operator, instead of the binary division operator, allows us to better control the constraints to avoid numerical issues. We set up a series of constraints on these operators that attempt to avoid circularly nested functions, i.e. a sqrt directly inside a square, or a log directly inside an exponential. Additionally we do not allow any ++ or - operators inside the inverse to avoid division by 0. In our testing we found this was necessary to ensure numerically stable equations and did not meaningfully alter the final accuracy of the equations.
We use a L2, or mean square error, loss function and a maximum equation complexity of 50. Complexity is measured as the total number of components in an equation; each constant, operator and variable contributes 1 complexity. We use the batching procedure within pysr with a batch size of 500. In our testing we find this was necessary to explore more equations more efficiently. We use 150 population and 2,000 cycles per iteration while setting weight_optimize parameter to $10^{-3}$ as suggested in the documentation to ensure the constants are optimized. We set the Parsimony and adaptive parsimony scaling to $10^{-8}$ and 500 respectively. These parameters control how much to punish high complexity solutions and we use a lower scaling value to encourage exploration of high complexity, but more accurate, equations. We perform the symbolic search for $2.5\times 10^{6}$ iterations, with the results summarized in Figure 2.
@@ 3.2 Selecting an Equation
For many applications, the goal of symbolic regression is to find a simple or interpretable equation. However, our aim is to find an accurate and computationally efficient equation, with little emphasis on readability or simplicity. To understand the properties of the final population of equations in this context, Figure 2 compares the computational efficiency against the accuracy of the final pool of equations. For accuracy we use the L2, or mean squared error, loss for a validation set of 1,000 $k,n$-$\mathcal{F}_{r}$ pairs that are drawn from the same distribution as is used for training. For computational efficiency we measure the execution time to render an 150 x 150 pixel image for a given set of parameters and calculate the $\chi^{2}$ log likelihood, assuming no covariance between pixels, with respect to a reference image. This is implemented in jax, same as pysersic, which allows for auto-differentiation capabilities. In this timing we also include the calculation of gradients of the log likelihood. This setup attempts to mimic the calculations used for sampling and optimization when these emulators are implemented into pysersic. This timing is repeated 500 times for each equation and we report the median. For comparison we repeat the timing procedure for the hybrid renderer, which serves as the default in pysersic and find an average execution time of 1.66 ms.
Figure 3: Visualization of the chosen equation compared to analytically or numerically calculated radial Fourier Transforms. We show four profiles, $n=0.5,1,2.5,4$, across two panels which show the profiles with logarithmic (top) and linear (bottom) scalings. In each panel we also show the linear residuals, the same for both panels but with different scaling of the $k$ axis. For $n=0.5$ and $n=1$, we show the analytically and numerically calculated Hankel transforms. For $n=2.5$ and $n=4$ the analytic forms are not known so we rely only on the numerically calculated transforms as the truth.
@@
Figure 4: Injection recovery tests to test if our chosen Sérsic Fourier Emulator, Eqn. 6 is accurate for profile fitting. We render a series of profiles using a heavily over-sampled pixel space renderer as the ‘truth‘ and fit each using a renderer based on the Fourier Emulator. The input values are compared to those recovered for all of the parameters of the profile. For the flux, since all of the input values are 100, we opt to show a histogram of the residuals. In general we find excellent agreement. For the position, ellipticity and positional angle we find recoveries that are accurate to a fraction of a percent. For $r_{\rm eff}$, the index $n$ and flux the average residual $0.5\%$ and the scatter $2\%$, calculated using the biweight location and scale respectively (Beers et al., 1990), for all three parameters. These results indicate that the rendering method utilizing the Sérsic Fourier emulator can be used for profile fitting with minimal loss in accuracy.
@@
In general we find there is an inverse correlation between accuracy and computational time. The more complex equations are generally more accurate but require more calculations and therefore have longer execution times. However even these most complex equations still have an execution time over twice as fast as the hybrid renderer. We find there is a plateau in this relationship. Above a complexity of about 30 the execution time continues to increase without any significant gains in the accuracy. In order to maximize both computational efficiency and accuracy we aim to select a point near the beginning of this plateau. Specifically we select the equation with the lowest execution time that has L2 loss $2\times 10^{-6}$. This yields the equation with a complexity of 39, which is displayed below, in Equation. 6.
@@
This equation is an outlier in the plane of execution time and accuracy that makes it particularly suited for our use case. Compared to equations of similar complexity in the final pool there are very few instances of the $k$ evaluation within exponential or square root evaluations. When rendering a profile the number of values of $k$ that need to be calculated scales with the number of pixels and these are computationally expensive operations. Given there is only one value of $n$ to evaluate for each profile the prevalence of this variable has less of an effect on the overall execution time. For the remainder of this paper we will use this equation used to emulate the radial Fourier transform of the Sérsic profile and denote it $\tilde{\mathcal{F}_{r}}$.
Our chosen equation is visualized and compared to the true radial Fourier transform in Figure 3. The emulated equation is shown at four different indices, $n=0.5,1,2.5$ and $4$, alongside the residuals when compared to the true value. For $n=0.5,1$, we include the analytically calculated radial Fourier transform alongside the numerically calculated one. The selected equation provides a good match to all the profiles across the entire $k$ range shown, with residuals constrained to $10^{-2}$. The scatter accuracy does depend on $n$, with the values of $n<1$ showcasing larger residuals. The emulated equation does diverge from the truth at high $k$ for these profiles as well but this is most evident in the top panel with logarithmic scaling. While both the chosen equation and the numerically calculated $\mathcal{F}_{r}$ trend to zero at high $k$, the behavior is slightly different. The emulator function declines quicker, at least for values of $n<1$.